A small business buys a computer for $3300. After 4 years the value of the computer is expected to be $200. For accounting purposes the business uses linear depreciation to assess the value of the computer at a given time. This means that if V is the value of the computer at time t, then a linear equation is used to relate V and t.
(a) Find a linear equation that relates V and t.

Answer by stanbon(75887)

A small business buys a computer for $3300. After 4 years the value of the computer is expected to be $200. For accounting purposes the business uses linear depreciation to assess the value of the computer at a given time. This means that if V is the value of the computer at time t, then a linear equation is used to relate V and t.
(a) Find a linear equation that relates V and t.
-----
You have 2 points: (0,3300) and (4,200)
----
slope = (200-3300)/(4-0) = -3100/4 = -775
-----
Form: V = mt + b
You know m = -775 and V = 3300 when t = 0
Solve for "b":
3300 = -775*0 + b
b = 3300
------
Equation:
V(t) = -775t + 3300
----------------------------
Cheers,
Stan H.
====================